A vertical scale on a spring balance reads from 0 to 250 N . The scale has a length of 10.5 cm from the 0 to 250 N reading. A fish hanging from the bottom of the spring oscillates vertically at a frequency of 2.75 Hz .

Ignoring the mass of the spring, what is the mass m of the fish?
Answer #1

A = 10.5cm = 0.105m

F = 250N

From F = KA

=) K = F/A = 250/0.105 = 21428.57N/m

Now for fish , f = 2.75Hz

Also f = w/2.pai = 1/2.pai × √k/m

=) (2.pai.f)^2 = k/m

=) m = k/(2.pai.f)^2 = 21428.57/(2×3.143×2.75)^2 = 71.7kg

=) Mass of fish m = 71.7kg